### v1

# MedPromptExtract (Medical Data Extraction Tool): Anonymization and Hi-fidelity Automated data extraction using NLP and prompt engineering

Abstract
A major roadblock in the seamless digitization of medical records remains the lack of interoperability of existing records. Extracting relevant medical information required for further treatment planning or even research is a time consuming labour intensive task involving the much valuable time of doctors. In this demo paper we present, MedPromptExtract an automated tool using a combination of semi supervised learning, large language models, natural language processing and prompt engineering to convert unstructured medical records to structured data which is amenable to further analysis.
1 Introduction
Extracting relevant data from non digitized medical records can be a challenging task, moreover in low and middle income countries(LMICs) such data may not be available in interoperable formats Bajpai and Wadhwa (2021). Additionally, data may be required in retrievable formats for research purposesLehne et al. (2019). Discharge summaries, which provide comprehensive information about a patient’s history, treatment, and post-discharge care instructions, are valuable sources of such data Komenan et al. (2023). Anonymizing such records is essential to maintain confidentiality when used for research Rothstein (2010). Moreover, the implementation of artificial intelligence (AI) and machine learning, along with the adoption of paperless workflows, is crucial for transforming healthcare data and improving clinical decision-making.
In this paper, we present MedPromptExtract a fully automated tool for effectively extracting data from discharge summaries while ensuring confidentiality.
2 Methodology
2.1 Dataset
This study was a part of our long term goal to identify predictors of Acute Kidney Injury (AKI) in patients admitted to Kokilaben Dhirubhai Ambani Hospital (KDAH), Mumbai, India. We collected 914 discharge summaries retrospectively from patients likely to have AKI based on serum creatinine values, a biochemical product which is marker for AKI as per KDIGOKhwaja (2012) 2012 and AACCEl-Khoury et al. (2021) 2020 guidelines. All discharge summaries followed a standard format, facilitating automated data extraction. Ethical clearance was obtained from the Institutional Ethical Committee (IEC) of KDAH, Mumbai (IEC - A Code: 037/2021).
2.2 Anonymisation of discharge summaries
We used EIGEN (Expert-Informed Joint Learning aGgrEatioN)Singh et al. (2023) for anonymization, leveraging large language models (LLMs) and semi-supervised learning techniques for high-fidelity information extraction from documents. We used Document Text Recognition (DocTR) Mindee (2021) to convert PDF files to JSON files for compatibility with EIGEN. EIGEN employs optical character recognition (OCR) and layout analysis to extract text and spatial structure from the discharge summaries. Labeling functions (LFs),which are are functions or heuristics that generate noisy or approximate labels for training data Ratner et al. (2016), were designed using positional heuristics to generate surrogate labels, significantly reducing annotation costs. Our training set consisted of 714 documents, of which only 20 were labeled, minimizing annotation costs.
Figure 1: Visualizing EIGEN
2.2.1 Model Implementation
We used LayoutLM Xu et al. (2020), a pre-trained deep neural network model, for information extraction from document images. The ”feature model” utilizes LayoutLM, while the ”graphical model” involves a probabilistic model that aggregates LF outputs to assign labels to unlabeled data. In our experiment, we had four classes to label: name, patient ID, location, and others. The batch size used was 32. The
Algorithm 1 Anonymisation of Discharge Summaries
Input: text triggers, word, bounding box
Parameters: Batch size, learning rate for both models
Output: Predicted Class Labels
1:  Getting surrogate labels: Define text triggers for LFs, define positional heuristics
2:  for Each Labelling Function do
3:     Match text triggers with words in the document
4:     if Text matched and bounding box satisfies positional heuristics then
5:        Assign class label to the word
6:     else
7:        Abstain
8:     end if
9:  end for
10:  Load data - Discharge Summaries, Set optimizer to AdamW
11:  for Epoch in range of number of epochs do
12:     Train feature model (LLM), graphical model
13:     return  Labels for all words for each document in the dataset
14:     Calculate Total Loss
15:     Back propagate losses and update model weights
16:  end for
17:  return Predicted Class Labels
learning rate for the feature model and graphical model was $5\times 10^{-5}$ and 0.01 respectively. The optimizer used was AdamW. We used predicted class labels to anonymize discharge summaries by removing patient names and locations, replacing patient IDs with serial numbers. This process was conducted within the medical records department of KDAH, ensuring complete confidentiality. Figure 1 shows the visualization of bounding boxes for class labels after implementing EIGEN.
2.3 Extracting text from anonymised records
2.3.1 Preprocessing: Using Stopwords
Stopwords, curated as keywords, guide the NLP model by excluding irrelevant text sections. Section headings, example ”Past History” and ”Discharge Date,” were used as stopwords.
Figure 2: Designed prompt
Algorithm 2 Extract Patient Information from anonymised Discharge Summaries
0:  Folder containing PDF files
0:  DataFrame containing extracted information
1:  Define stop_keywords as a list of keywords to halt text extraction
2:  Set folder_path to the directory containing the PDF files
3:  Initialize an empty DataFrame df with specified columns
4:  for each filename in the folder_path do
5:     if the filename ends with ”.pdf” then
6:        Construct the full file path
7:        Initialize a dictionary data_dict to store extracted data
8:        Initialize an empty string extracted_text
9:        for each page in the PDF do
10:           Extract text from the page, Append the extracted text to extracted_text
11:        end for
12:        Replace newline characters in extracted_text with spaces
13:        Use regular expressions to extract data based on predefined patterns and stop_keywords
14:        Store extracted information in data_dict
15:        Normalize data_dict to ensure all lists have the same length
16:        Append data from data_dict to the DataFrame df
17:        Print any exceptions encountered
18:     end if
19:  end for
20:  return  the DataFrame df
Figure 3: Flowchart of anonymisation and data extraction process
Feature Accuracy Sensitivity Specificity Precision
F1
Score
AUC
AKI Mentioned 0.96 0.5 1 1 0.67 0.75
Angiography Done 0.98 1 0.98 0.86 0.92 0.98
Consultation by Nephrologist 0.81 0.12 0.95 0.33 0.18 0.53
Diuretic Given 0.98 1 0.97 0.92 0.96 0.98
Fluid Restriction Advised 0.98 1 0.98 0.67 0.8 0.98
General Anaesthesia 0.81 0.65 0.9 0.79 0.71 0.77
Hyperkalemia Not computed as there were no true positives
Hypernatremia Not computed as there were no true positives
Hypertension 0.98 1 0.98 0.75 0.86 0.98
Hypokalemia Not computed as there were no true positives
Hyponatremia Not computed as there were no true positives
ICU Admission 0.71 0.5 0.95 0.93 0.65 0.72
Imaging Procedure with Contrast 0.94 0.88 0.95 0.78 0.83 0.9
Oxygen Saturation Drop 0.96 0.5 1 1 0.67 0.75
Tachycardia 1 1 1 1 1 1
Ventilator Used 0.98 1 0.98 0.8 0.89 0.98
Table 1: Validation of model responses with clinicians responses
2.3.2 NLP Model Architecture
  • Regular Expressions: Employed to match specific patterns within the text, facilitating information extraction.
  • Stopword Filtering: Ensures exclusion of irrelevant text sections related to patient identification.
The extracted information was organized and stored in a structured DataFrame, with each stopword serving as a feature label for information following it.
2.4 Utilizing Prompt Engineering on Gemini to extract information
In this step, we analyzed free-flowing text under the heading ”course in hospital” to identify predictors of AKI. Nineteen features were identified through domain knowledge for extraction using the generative ability of LLMs and prompt engineering. A single prompt(Figure 2) was used to extract all clinical impressions, utilizing the Gemini APIPichai (2023). The API’s generative model, gemini-pro, was configured with temperature parameter set to 0 to ensure precise and deterministic outputs. Two iterations of model responses were subjected to inter-rater agreement using Kappa coefficient.
2.5 Validation with human annotations
Data extracted using LLM was validated against human annotations for 48 discharge summaries. Two clinicians annotated all summaries, and inter-rater agreement was assessed using the Kappa coefficient. Accuracy, precision, recall, F1 score and AUC were calculated using the ScikitLearn libraryPedregosa et al. (2011) in Python. Figure 3 illustrates the anonymisation and data extraction process of MedPromptExtract
3 Results
Out of the 914 discharge summaries processed, four were excluded initially due to format errors in the source PDFs received from the hospital. This anomaly was communicated to the hospital’s internal quality control cell. Upon running the prompt, it was observed that 58 summaries had an empty ”course in hospital” field, with two of them also included in the randomly selected annotations for human validation. Ultimately, outputs were obtained for 852 summaries, with 48 used for human validation. During annotation, it was found that seven of the 19 prompt questions had negative responses for all instances that is no true positives, leading to their exclusion from the final analysis.
Annotating a single summary took an average of 12 minutes, totaling 9.6 man-hours for 48 annotations. The compute time required for one iteration of 853 summaries was 100 minutes, equating to nearly 7 seconds per summary. Gemini’s API operates independently of the user’s local hardware, with computations conducted on their servers.
Inter-rater agreement between the two doctors was over 0.8, and within the two model iterations, it was 1. The validation of model responses against human responses is detailed in table 1 , where AUC for 7 of the 12 features was above 0.9, indicating high fidelity of the data extraction process.
4 Limitations
Our pipeline is custom-developed for the discharge summaries of the source hospital. While generalizable, it would require minor adjustments to the training of EIGEN and the prompts to be used for other similar usecases.
5 Conclusion
Translating machine learning tools to real life hospital scenarios is a challenge given the practical constraints set by data quality and availability. This demo paper brings forth the ability of our model to perform automated anonymisation and data extraction with hi-fidelity utilizing very less labelled data, significantly reducing anonymisation costs, thus providing a good framework for digitization and further analysis of healthcare data. This method makes it possible to make data available in readily consumable format while maintaining anonymity. MedPromptExtract will be subsequently deployed in the source hospital and will be integrated with their EHR system for downstream applications.
6 Url for video